OpenStudy (anonymous):

4 times a number is 6 more than the number. what is the number?

OpenStudy (anonymous):

4x=6+x 4x-x=6 3x=6 x=6/3 x=2

OpenStudy (anonymous):

The number is 2

OpenStudy (anonymous):

You can translate word problems by carefully looking at the words. \[4 \rightarrow 4\]\[times \rightarrow \times\]\[a \space number \rightarrow x\]\[is \rightarrow =\]\[\text{6 more than} \rightarrow + 6\]\[the \space number \rightarrow x\]Putting it all together: \(4 \times x = 6 + x\) which is the same as \(4x = x + 6\)which @monokerous has up there.